26. Suppose A is 2 x 2, and detA= -1, tr(A) = 0. Show that A² = 1.

Suppose A is 2 × 2, and detA = −1, tr(A) = 0. Show that A2 = I
Answer #1

ginen olet (A) = -) and tr(A)=0 so ad-cb=-1 and a+d=0 > d=-a ad-cb=-1 -a²-c6==1= a + b c = 1 Now A= (aby) - 18tbc = (ca-ca ab